# Map2Text: New Content Generation from Low-Dimensional Visualizations

### v1

@@ Abstract.
Low-dimensional visualizations, or ”projection maps” of datasets are widely used across scientific research and creative industries as effective tools for interpreting large-scale and complex information. These visualizations not only support understanding existing knowledge spaces but are often used implicitly to guide exploration into unknown areas. While powerful methods like TSNE or UMAP can create such visual maps, there is currently no systematic way to leverage them for generating new content. To bridge this gap, we introduce Map2Text, a novel task that translates spatial coordinates within low-dimensional visualizations into new, coherent, and accurately aligned textual content. This allows users to explore and navigate undiscovered information embedded in these spatial layouts interactively and intuitively. To evaluate the performance of Map2Text methods, we propose Atometric, an evaluation metric that provides a granular assessment of logical coherence and alignment of the atomic statements in the generated texts. Experiments conducted across various datasets demonstrate the versatility of Map2Text in generating scientific research hypotheses, crafting synthetic personas, and devising strategies for testing large language models. Our findings highlight the potential of Map2Text to unlock new pathways for interacting with and navigating large-scale textual datasets, offering a novel framework for spatially guided content generation and discovery.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@
Illustration of Map2Text task on crafting synthetic persona. Refer to caption
Figure 1. Illustration of a Map2Text task which generates new content based on a 2D projection map. The figure provides an example of generating new red-teaming strategies for testing large language models. Given a low-dimensional visualization precomputed from certain text corpora and a query position (green dot), Map2Text challenges a model to generate new, coherent, and meaningful textual content that would correspond to that location in the visualization.
@@ 1. Introduction
Low-dimensional visualizations, or “projection” maps, are powerful tools that help users understand large textual corpora, much like maps aid in navigating physical landscapes. Typically, these visualizations are created by first generating high-dimensional embeddings using text embedding models (e.g.,  (Muennighoff et al., 2022)), followed by dimension reduction algorithms such as TSNE (Van der Maaten and Hinton, 2008), LargeVis (Tang et al., 2016), or UMAP (McInnes et al., 2018). This process results in a 2D (or 3D) map that preserves semantic relationships between texts, enabling users to intuitively explore complex datasets and identify interesting patterns (Low et al., 2020; Noichl, 2021; Yang et al., 2022; Tschisgale et al., 2023; Wang et al., 2022). This utility raises an intriguing question: much like how people use geographical maps to explore new territories, can we use these visualizations not only to navigate existing knowledge but also to generate new content?
This is the central idea behind our proposed task, Map2Text. More specifically, Map2Text aims to generate textual content corresponding to a given position within a low-dimensional visualization map. The fundamental assumption that Map2Text relies on is that semantically similar texts are positioned close together on a low-dimensional visualization, creating a “map” that conveys rich contextual information beyond the individual texts themselves. By leveraging this spatial context, Map2Text aims to guide the generation of new, coherent, and meaningful content.
@@
Many downstream applications could benefit from Map2Text. In scientific idea generation (Gu and Krenn, 2024; Zhou et al., 2024; Wang et al., 2023), researchers can select positions on the visualization map where concepts are converging or underexplored. By generating text corresponding to these research gaps, Map2Text can inspire novel research directions and foster innovation. Another promising application is to test the safety and robustness of systems, including large language models (LLMs), which continuously benefits from new adversarial tests (Ganguli et al., 2022; Perez et al., 2022). Typically, developing new red teaming strategies relies on expert knowledge, which requires intensive learning processes and is not scalable. With Map2Text, security experts could navigate visualizations of existing adversarial attempts and generate new strategies at scale to identify and mitigate vulnerabilities in LLMs. Overall, Map2Text could be a valuable tool for any field that relies on continuous innovation and new content generation, through enabling the exploration into uncovered regions in an information space.
@@
The Map2Text task presents several significant challenges. First, the process of reducing high-dimensional text embeddings to a two-dimensional map often introduce distortions and ambiguities in spatial representation. These artifacts may lead to inaccuracies when mapping back from a 2D point to the high-dimensional semantic space, affecting the relevance and coherence of the generated text (Rudin et al., 2022). Second, modeling text generation conditional on specific locations in the embedding space is non-trivial. Developing models that effectively integrate spatial information into the text generation process is essential to ensure that the output aligns with the intended semantics (Hu et al., 2017). Third, evaluating the quality and relevance of the generated content poses a challenge. Traditional evaluation metrics may not adequately capture the alignment between the generated text and its intended position on the map, necessitating the development of new evaluation frameworks (Goyal et al., 2022). Fourth, scalability and computational resource constraints are significant considerations. Handling large textual corpora and generating real-time responses based on user-selected points on the map can be computationally intensive, requiring optimization of algorithms for efficiency without compromising accuracy. Addressing these challenges is critical for the successful implementation and practical application of the Map2Text task.
Along this direction, we make the following initial contributions:
@@
  • We introduce Map2Text, a novel task of generating text given positions on a visualization map, transforming existing visualization maps into tools for new content creation [^1] Code to replicate our experiments can be found at the following anonymous repo: https://anonymous.4open.science/r/map2text-27FD.
  • We propose Atometric, a reasoning-based evaluation metric that uses atomic statements to evaluate text generation quality at multiple levels of strictness [^2] https://anonymous.4open.science/r/atometric-17A3/.
  • We demonstrate Map2Text’s potential on four datasets, benchmarking the performance of seven candidate models and human baselines. We obtain interesting findings and showcase the potential of Map2Text across domains.
@@
The remaining sections are organized as follows: Section 2 defines the problem formulation of Map2Text. Section 3 introduces the proposed evaluation metric, Atometric . Section 4 presents the experimental setup and results across multiple datasets. Section 5 reviews related work, and finally, Section 6 discusses some limitations of the paper and outlines future research directions.
Figure 2. An illustrative example of Atometric on generating new red teaming strategies for LLMs. Atometric decomposes the generated text into a set of atomic statements, which are individually compared against the reference text to evaluate how well each statement is supported under different levels of strictness—thereby measuring the “precision” of the generated text. Conversely, the reference text can be decomposed and compared against the generated text to assess “recall”. Both decomposition and verification processes are automated using LLMs with a prompt template, detailed in Appendix B.
2. Map2Text: A New Task
In this section, we introduce the formulation of Map2Text and its evaluation framework.
@@ 2.1. Problem Definition
Let ${\mathcal{V}}$ represent a low-dimensional projection of a text corpus $\mathcal{D}=\{s_{1},s_{2},\ldots,s_{N}\}$, where $s_{i}\in{\mathcal{S}}$ is the $i$-th text entry and ${\mathcal{S}}$ denotes the set of all possible text strings. For each text entry $s_{i}$, ${\mathcal{V}}$ includes a corresponding low-dimensional (e.g., 2D) position ${\bm{x}}_{i}\in{\mathbb{R}}^{2}$, forming a set of positions $\mathcal{X}=\{{\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{N}\}$. The visualization ${\mathcal{V}}$ is typically derived from $\mathcal{D}$ using a mapping function $f:{\mathcal{S}}\to{\mathbb{R}}^{2}$ that jointly maps each text entry $s_{i}$ to its position ${\bm{x}}_{i}$. This mapping function $f$ typically integrates a high-dimensional text embedding model and a dimensionality reduction algorithm. However, in many practical situations, the visualizations are precomputed and the details of the mapping function $f$ are unavailable to the users. Even when $f$ is known, reconstructing the visualization pipeline from scratch can be computationally expensive and subject to randomness. Thus, we assume no access to $f$ but rely on the assumption of locality preservation—points close to each other in the low-dimensional space $\mathcal{X}$ correspond to semantically similar text in the high-dimensional space ${\mathcal{S}}$. This allows us to use proximity in the 2D projection as a guide for generating contextually relevant text without requiring the original mapping function $f$.
Definition
The Map2Text task is defined as follows: given a query position ${\bm{x}}_{q}\in\mathbb{R}^{2}$ that has not been occupied by an existing text entry, a model $g$ is tasked with generating a text entry $\hat{s}$ that corresponds to ${\bm{x}}_{q}$. Specifically, the generated text $\hat{s}$ should be semantically and logically aligned with the ground truth text $s\in{\mathcal{S}}$. Here, a ground truth $s$ is a text entry that would have been mapped to ${\bm{x}}_{q}$ by the original mapping function $f$. Thus, the objective of the Map2Text task is to effectively learn an inverse mapping $g=\widehat{f^{-1}}:{\mathbb{R}}^{2}\to{\mathcal{S}}$ that reconstructs the relationship between the provided visualization and text entries.
@@ 2.2. Potential Solutions
Given the typical pipeline of $f$, there are multiple natural designs, each offering a distinct process on how to reconstruct text entries from low dimensional space.
@@
The most straightforward approach is to directly learn an inverse mapping from the 2D positions to the corresponding text entries, requiring the model to memorize the entire visualization map, associate each position with its corresponding text, and generalize the learned relationship to unoccupied positions. This could be achieved by fine-tuning a pre-trained language model that takes the input of a 2D coordinate and outputs a text string. This process can be described as $g:{\mathbb{R}}^{2}\to{\mathcal{S}}$.
If high-dimensional embeddings of the text corpus are available, we could introduce an intermediate step by mapping the 2D positions to a high-dimensional textual embedding space before reconstructing the text. This can be achieved using a neural network. The embedding is then decoded back into text using an embedding inversion model such as the vec2text model (Morris et al., 2023) (note that an embedding inversion model requires access to the original embedding model $f_{e}:{\mathcal{S}}\to{\mathbb{R}}^{H}$). This approach can be expressed as $g:{\mathbb{R}}^{2}\to{\mathbb{R}}^{H}\to{\mathcal{S}}$, where ${\mathbb{R}}^{H}$ denotes the $\underline{H}$igh-dimensional embedding space.
Another approach may first generate an initial textual description or of the 2D position. This description is then be used to generate the final text, which can be done through prompting a language model. In this case, the inverse mapping is decomposed into two stages: first, the model constructs a preliminary text prompt, then uses this prompt to generate the final output, which could be described as $g:{\mathbb{R}}^{2}\to{\mathcal{S}}\to{\mathcal{S}}$.
A more detailed description and analysis of these methods and their performance is provided in Section 4.
2.3. Evaluation Framework
Like other tasks of new content generation, evaluating the Map2Text task presents unique challenges due to the lack of ground-truth content at arbitrary, previously unexplored positions. While domain experts could compare the quality and accuracy of generated contents given specific query positions, this online process is not scalable and thus unsuitable for automatic benchmarking.
To overcome this, we adopt an offline benchmarking approach by partitioning a precomputed visualization ${\mathcal{V}}$ into a training set and a testing set. In this approach, the training set includes both original text entries and their corresponding positions in the 2D map. This allows the model to understand the relationship between the text and positions. In contrast, for the testing set, only the positions are made available to the model, serving as query points for which the model generates the associated text. The generated text is then compared against the ground-truth text held out from the original visualization, which serves as a gold standard reference.
This framework addresses the scalability issue of human judgment but faces another challenge in aligning the generated content with the reference text. For instance, the generated text might use different wording or phrasing while still conveying the same underlying idea as in the reference. Traditional lexical-level metrics may fail to capture such equivalence, penalizing otherwise valid outputs. This motivates the proposal of Atometric, an evaluation metric designed to assess semantic and logical alignment between the generated and reference texts across varying levels of strictness.
3. Atometric: A New Evaluation Metric
A key challenge of evaluating Map2Text is to identify an appropriate metric that accurately assesses the alignment between the generated content $\hat{s}$ and the gold standard reference $s$ in our specific context. We outline the essential requirements for such a metric and introduce a novel metric tailored to meet these criteria.
3.1. Motivation
Unlike traditional content generation tasks, which aims to replicate ground-truth texts, Map2Text requires evaluating whether the generated content is capable of inspiring the user to explore the map in a similar way—despite how much the surface expressions match or differ, the generated content must accurately reflect the semantic relationships and logical structures of reference content.
Existing evaluation metrics struggle to capture the semantic and logical complexity required for the Map2Text task. Traditional metrics such as ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002) rely on lexical-level n-gram overlaps, penalizing legitimate variations in phrasing and structure even when the generated content accurately conveys the same meaning. Recent metrics like BERTScore (Zhang et al., 2019) and BLEURT (Sellam et al., 2020) address some of these limitations by using contextual embeddings to measure semantic similarity. However, these metrics lack interpretability, as they offer a high-level similarity scores without explaining the nuances of why certain outputs are deemed similar or dissimilar. Moreover, these metrics evaluate holistic sentence- or paragraph-level similarity, neglecting the internal coherence of fine-grained statements. This limitation is crucial for Map2Text , where maintaining the precise relationships and logical structures between entities is critical.
3.2. Features
To address the limitations of existing metrics, we propose Atometric, a novel metric specifically designed to evaluate the semantic and logical coherence of generated content in the Map2Text task. An illustrative example of Atometric is shown in Figure 2. Atometric has the following features:
Atomic Statement Alignment.
Inspired by FActScore (Min et al., 2023), Atometric focuses on the alignment of atomic statements—the fundamental units of meaning that express relationships between entities and actions within a text. By directly evaluating these atomic statements instead of the complex full texts or fragmented ngrams, Atometric assesses whether the generated content preserves the logical relationships and semantic properties present in the reference.
Multiple Levels of Strictness.
Atometric introduces multiple strictness levels to capture nuanced levels of alignment. Inspired by definitions in textual entailment (Bowman et al., 2015a; Sadat and Caragea, 2022; Williams et al., 2018), these strictness levels offer varying degrees of relevance, from direct support to more relaxed criteria such as topic relevance. This flexibility ensures that the metric can accommodate the inherent variability of natural language while remaining faithful to the core semantic content.
Comprehensive Evaluation Metrics:
Atometric provides precision, recall, and F1-score measurements that jointly evaluate the correctness and completeness of the generated content. Precision measures the logical consistency and accuracy of generated statements, while recall captures the extent to which the generated content covers all key statements of the reference. This approach ensures a thorough assessment of both semantic accuracy and content coverage.
3.3. Definition
The formal definition of Atometric is as follows. Given a pair of generated and reference text $(\hat{s},s)$, We first extracts a set of atomic statements from $\hat{s}$ and $s$ using an LLM, denoted as ${\mathcal{A}}_{\hat{s}}$ and ${\mathcal{A}}_{s}$, respectively. The two sets of atomic statements are further verified by another LLM at multiple levels of strictness $l$ to evaluate both precision and recall. Mathematically, Atometric precision, recall, and $F_{1}$-score at level $l$ is defined as:
(1) $\text{Atometric}_{P}$ $=\dfrac{1}{|{\mathcal{A}}_{\hat{s}}|}\sum_{a\in{\mathcal{A}}_{\hat{s}}}{\mathbb{I}}^{l}[a\,\text{is relevant to}\,s]$
(2) $\text{Atometric}_{R}$ $=\dfrac{1}{|{\mathcal{A}}_{s}|}\sum_{a\in{\mathcal{A}}_{s}}{\mathbb{I}}^{l}[a\,\text{is relevant to}\,\hat{s}]$
(3) $\text{Atometric}_{F1}$ $=\dfrac{2\times\text{Atometric}_{P}\times\text{Atometric}_{R}}{\text{Atometric}_{P}+\text{Atometric}_{R}}$
where ${\mathbb{I}}^{l}[\cdot]$ is an indicator function at strictness level $l$ that checks if an atomic statement is relevant to the corresponding reference or generated text in the sense of textual entailment (Sadat and Caragea, 2022; Bowman et al., 2015a; Williams et al., 2018). We show an overview of criteria in Table 1 and provide a detailed template in Appendix Figure 5 that could be easily adapted to different datasets with minimal change.

### v2

# MapExplorer: New Content Generation from Low-Dimensional Visualizations

@@ Abstract.
Low-dimensional visualizations, or ”projection maps” are widely used in scientific research and creative industries to interpret large-scale and complex datasets. These visualizations not only support the understanding of existing knowledge spaces but are often used implicitly to guide exploration into unknown areas. While such visualizations can be created through various methods such as TSNE or UMAP, there is no systematic way to leverage them for generating new content. To bridge this gap, we introduce MapExplorer, a novel knowledge discovery task that translates coordinates within any projection map into new, coherent, and accurately aligned textual content. This enables users to interactively explore and discover insights embedded in these maps. To evaluate the performance of MapExplorer methods, we propose Atometric, a granular metric inspired by ROUGE, which quantifies logical coherence and statement alignment of generated text against references. Experiments across diverse datasets demonstrate the versatility of MapExplorer in generating scientific research hypotheses, crafting synthetic personas, and devising strategies for attacking large language models, even with straightforward baseline methods. By bridging visualization and generation, our findings highlight the potential of MapExplorer to unlock new pathways of intuitive human-AI collaboration for large-scale dataset exploration [^1] Codebase: (1) https://github.com/xingjian-zhang/map2text; (2) https://github.com/xingjian-zhang/atometric..
@@
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1454-2/25/08doi: 10.1145/XXXXXX.XXXXXXccs: Computing methodologies Spatial and physical reasoningccs: Computing methodologies Natural language generationccs: Human-centered computing Visualization toolkitsccs: Human-centered computing Empirical studies in visualizationccs: Human-centered computing Scientific visualizationccs: Human-centered computing Information visualization
@@
Illustration of MapExplorer task on crafting synthetic persona. Refer to caption
[…]
@@ 1. Introduction
Low-dimensional visualizations, or “projection maps”, are powerful tools that help users understand large textual corpora, much like maps aid in navigating physical landscapes. These maps are often constructed by various recipes. For example, researchers first generate text embeddings and subsequently apply dimension reduction techniques such as t-SNE (Van der Maaten and Hinton, 2008), LargeVis (Tang et al., 2016), or UMAP (McInnes et al., 2018). Alternatively, these maps can be constructed from structural relationships within the data, including explicit relations such as citations or implicit relations such as co-occurrences, where the spatial positioning of elements reflects the underlying network topology. Regardless of the construction method, these maps aim to preserve meaningful relationships between textual or other types of elements, enabling users to intuitively explore complex datasets and discover interesting patterns (Low et al., 2020; Noichl, 2021; Yang et al., 2022; Tschisgale et al., 2023; Wang et al., 2022). This utility raises an intriguing question: much like how people use geographical maps to explore new territories, can we use these projection maps not only to navigate the existing content space but also to explore new knowledge?
This is the central idea behind our proposed novel knowledge discovery task, MapExplorer. As illustrated in Figure 1, MapExplorer aims to generate textual content corresponding to a given position within a low-dimensional visualization map. This effectively converts the static visualization into an interactive and generative tool that facilitates exploration.
MapExplorer is built on two principles: (1) Semantically similar texts are positioned close together on a low-dimensional visualization, creating a map that conveys rich contextual information beyond the individual texts themselves. Leveraging this spatial context, MapExplorer aims to guide the generation of new, coherent, and meaningful content. (2) Any candidate method should be visualization-agnostic, operating on the final map without requiring access to or knowledge of the underlying visualization process. This design choice ensures broad applicability across different visualization techniques and use cases, as visualization maps are usually provided as the final products in most practical scenarios (Map of Science, [n. d.]; MedViz, [n. d.]).
@@
Many downstream applications could benefit from MapExplorer. In scientific idea generation (Gu and Krenn, 2024; Zhou et al., 2024; Wang et al., 2023), researchers can select positions on the visualization map where concepts are converging or underexplored. By generating potential research ideas or hypotheses corresponding to these research gaps, MapExplorer can inspire novel research directions. Another promising application is to test the safety of large language models (i.e. red teaming), which continuously benefit from new adversarial tests (Ganguli et al., 2022; Perez et al., 2022). Typically, developing new red-teaming strategies relies heavily on expert knowledge. With MapExplorer, security experts could navigate visualizations of existing adversarial attempts and generate new strategies at scale to identify vulnerabilities in LLMs. Overall, MapExplorer could be a valuable tool for any field that relies on continuous innovation, through enabling the exploration into uncovered regions in an information space.
@@
The MapExplorer task presents several significant challenges. First, the visualization process often introduces distortions in spatial representation. These artifacts may lead to inaccuracies when mapping back from a 2D point to textual contents, affecting the relevance of the generated text (Rudin et al., 2022). Second, modeling text generation conditional on specific locations in the visualization map is non-trivial. Developing models that effectively integrate spatial information into the text generation process is essential to ensure that the output aligns with the intended semantics (Hu et al., 2017). Third, evaluating the quality and relevance of the generated content poses a challenge. Traditional evaluation metrics may not adequately capture the alignment between the generated text and its intended position on the map, necessitating the development of new evaluation frameworks (Goyal et al., 2022).
Contributions
We highlight the following initial contributions:
@@
  • We introduce MapExplorer, a novel task of generating text given positions on a visualization map, transforming existing visualization maps into tools for new content creation.
  • We develop a comprehensive evaluation framework for MapExplorer that enables offline assessment of generated content at scale.
  • We demonstrate MapExplorer’s potential on four datasets using a series of straightforward baseline methods. Our findings highlight MapExplorer’s versatility across various domains.
Disclaimer
The primary goal of this study is to introduce and demonstrate the feasibility of the MapExplorer task using various techniques. We do not focus on optimizing or endorsing specific methods for MapExplorer. While this study emphasizes quantitative and scalable offline evaluation, our tasks and methods are also prepared for potential online evaluation procedures, such as those described by (Si et al., 2024). For those interested, a live and anonymous demo of MapExplorer is available for ad hoc exploration at https://mapexplorer-app.streamlit.app/.
@@
This paper is structured as follows: Section 2 outlines the novel problem formulation of MapExplorer. In Section 3, we introduce the proposed evaluation framework and highlight the need for a new evaluation metric. Section 4 details the experimental setup and presents results across multiple datasets. Section 5 provides a review of related work. Finally, Section 6 discusses the paper’s limitations and suggests directions for future research.
[…]
2. MapExplorer: A New Task
@@ 2.1. Problem Definition
Consider a text corpus $\mathcal{D}=\{s_{1},s_{2},\ldots,s_{N}\}$, where each $s_{i}$ represents a textual entry. A projection map is defined as the set ${\mathcal{V}}=\{({\bm{x}}_{i},s_{i})\mid s_{i}\in\mathcal{D},{\bm{x}}_{i}\in\mathbb{R}^{2}\}$, where each text entry $s_{i}$ is associated with a corresponding position ${\bm{x}}_{i}$ in a low-dimensional space, i.e., a 2D space in most existing visualizations.
The MapExplorer task is defined as follows: given a query position ${\bm{x}}_{q}$ on the final visualization map—one not previously occupied by an existing text entry—MapExplorer aims to generate a text entry $\hat{s}$ that would have been mapped to ${\bm{x}}_{q}$.
For the generality and practicality of the task, this process operates on the map ${\mathcal{V}}$ as provided, but without requiring access to, or the ability to re-compute the original visualization procedures. This “visualization-agnostic” assumption aligns with the common scenario in which end-users only have access to the final low-dimensional visualization rather than the entire visualization pipeline or knowledge about the visualization method employed.
@@ 2.2. Potential Solutions
Given the problem definition, we discuss three natural designs on potential solutions here. Each of them offers a distinct process on how to construct textual content from the low-dimensional space. Detailed implementation of these solutions is provided in Section 4.
Direct Mapping.
@@
A straightforward approach is to learn an inverse mapping from 2D coordinates to textual entries. The model essentially ”memorizes” the map by associating each point with the corresponding text entry and generalizing to positions not currently occupied. In practice, this can be achieved by fine-tuning a pre-trained language model that takes a 2D coordinate as input and outputs a text string.
Intermediate High-Dimensional Embedding.
Another strategy involves converting the 2D coordinates back into a higher-dimensional text embedding space before generating text. This inverse mapping is divided into two steps: first, a neural network projects the 2D coordinates into a specific textual embedding space; second, an embedding-inversion model, such as vec2text (Morris et al., 2023), decodes those embeddings into text.
Two-Stage Text Generation.
An alternative approach is to generate a concise textual ”descriptor” for the 2D coordinate and then feed this descriptor into a language model to produce the final text. This method splits the mapping from a 2D coordinate to text into two steps: first, creating an initial prompt that summarizes or contextualizes the coordinate; second, using that prompt to obtain the final text output.
3. Evaluation
3.1. Offline Framework
Evaluating the MapExplorer task poses unique challenges similar to other new-content-generation problems. A major obstacle is the lack of ground-truth text at arbitrary, previously unexplored positions on the map. Although domain experts can judge the quality and relevance of generated outputs online (e.g., reviewing them one by one), such manual assessments are usually subjective and do not scale (Si et al., 2024), making them impractical for large-scale benchmarking.
To address these issues, we adopt an offline approach. Starting with a visualization ${\mathcal{V}}$, we split it into a training set and a held-out testing set. The training set provides both original text entries and their 2D positions, allowing a model to learn how textual content relates to different coordinates. In contrast, we withhold the text in the testing set, leaving only the 2D positions. During evaluation, the model’s generated texts are compared to the ”gold standard” references. Beyond objectivity and scalability, another advantage of this reference-based evaluation is that it eases the burden of evaluating the novelty of generated text, which is particularly challenging and often requires substantial domain knowledge from human experts (Si et al., 2024). As long as the gold-standard references are considered novel with regard to the training set, generated text that closely aligns with the references should also be novel.
While this setup reduces the reliance on human judgment, it also highlights another challenge: existing metrics, whether focused on lexical overlaps (e.g., BLEU and ROUGE) or embedding-based similarity (e.g., BERTScore and BLEURT), often struggle to handle valid paraphrases or to pinpoint which essential ideas have been omitted. In other words, a desirable generated text could convey similar novel ideas to the reference but have used a different lexical presentation or granularity of details. This limitation calls for a specialized evaluation method that can accommodate flexible phrasing yet still verify whether the generated text aligns with all key elements of the reference. In this work, we propose a new metric to address this challenge. We refer the readers to Section 6 for additional limitations of this offline evaluation framework.
3.2. Motivation for a New Metric
In MapExplorer, each generated text should ideally be both correct (i.e., free of logical or factual inconsistencies) and complete (i.e., capturing necessary core ideas from references) with regard to a gold-standard reference. Standard evaluation metrics, however, often struggle to balance these two objectives.
Classical approaches such as ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002) approximate precision and recall via n-gram overlaps. Though this conceptually aligns with the goals of accuracy and completeness, their reliance on strict lexical matching can penalize acceptable paraphrases. More recent embedding-based methods (e.g., BERTScore (Zhang et al., 2019) or BLEURT (Sellam et al., 2020)) offer more flexibility for reworded text but usually produce only a single similarity score. Such an aggregate score does not reveal precisely which critical ideas are included or whether any unintended statements have been introduced.
3.3. Atometric: A New Evaluation Metric
We now propose Atometric, a novel metric designed to address the shortcomings discussed above—particularly the need for fine-grained evaluation that balances correctness and completeness with regard to a reference. By focusing on atomic statements instead of entire sentences or overlapping n-grams, Atometric captures how well each key piece of information from the reference is preserved in the generated text. Figure 2 provides an illustrative overview.
Formal Definition
Let $\hat{s}$ be the generated text and $s$ the reference text. We first use an LLM to extract two sets of atomic statements, ${\mathcal{A}}_{\hat{s}}$ and ${\mathcal{A}}_{s}$. At a chosen strictness level $l$, another LLM checks how each statement in ${\mathcal{A}}_{\hat{s}}$ is entailed by $s$, and vice versa. Formally, Atometric precision ($\text{{Atometric}{}}_{P}$), recall ($\text{{Atometric}{}}_{R}$), and F1-score ($\text{{Atometric}{}}_{F1}$) at level $l$ are:
(1) $\text{{Atometric}{}}_{P}$ $=\frac{1}{|{\mathcal{A}}_{\hat{s}}|}\sum_{a\in{\mathcal{A}}_{\hat{s}}}{\mathbb{I}}[s\,\text{entails}\,a\text{ at level }l],$
(2) $\text{{Atometric}{}}_{R}$ $=\frac{1}{|{\mathcal{A}}_{s}|}\sum_{a\in{\mathcal{A}}_{s}}{\mathbb{I}}[\hat{s}\,\text{entails}\,a\text{ at level }l],$
(3) $\text{{Atometric}{}}_{F1}$ $=\frac{2\times\text{{Atometric}{}}_{P}\times\text{{Atometric}{}}_{R}}{\text{{Atometric}{}}_{P}+\text{{Atometric}{}}_{R}},$
where ${\mathbb{I}}[\cdot]$ is an indicator function induced by an LLM verifier. Table 1 shows an overview of the criteria for each strictness level, and Appendix C provides details on the LLM prompt template.
Relation to ROUGE
Atometric generalizes ROUGE in two important ways. First, it replaces n-gram overlaps with atomic statements, providing an appropriately-grained view of which specific facts or relationships are captured. Second, it relaxes exact lexical matching in favor of natural language entailment, allowing legitimate paraphrases and rewordings to count as valid matches.
In summary, the proposed metric has the following features:
  • Atomic Statement Alignment. Similar to FActScore (Min et al., 2023), Atometric breaks both reference and generated texts into atomic statements (i.e., factoids). Evaluating these statements ensures the comparison is conducted at the right granularity of ideas and avoids penalizing valid paraphrases while ensuring the critical ideas are captured.
  • Multiple Levels of Strictness. Inspired by textual entailment (Bowman et al., 2015a; Sadat and Caragea, 2022; Williams et al., 2018), Atometric uses a range of strictness levels in determining entailment to atomic statements. Looser levels check for broad consistency; stricter ones require explicit or paraphrased matches. This approach accommodates linguistic variety while preserving logical integrity.
  • Precision, Recall, and F1. Atometric measures correctness via precision (the fraction of generated atomic statements entailed by the reference) and completeness via recall (the fraction of reference atomic statements entailed by the generation). F1-score combines these two into a comprehensive metric.